# Designing Robots to Help Women*

### v1

@@ Abstract
Robots are being designed to help people in an increasing variety of settings–but seemingly little attention has been given so far to the specific needs of women, who represent roughly half of the world’s population but are highly underrepresented in robotics. Here we used a speculative prototyping approach to explore this expansive design space: First, we identified some potential challenges of interest, including crimes and illnesses that disproportionately affect women, as well as potential opportunities for designers, which were visualized in five sketches. Then, one of the sketched scenarios was further explored by developing a prototype, of a robotic helper drone equipped with computer vision to detect hidden cameras that could be used to spy on women. While object detection introduced some errors, hidden cameras were identified with a reasonable accuracy of 80% (Intersection over Union (IoU) score: 0.40). Our aim is that the identified challenges and opportunities could help spark discussion and inspire designers, toward realizing a safer, more inclusive future through responsible use of technology.
@@ I INTRODUCTION
Within the general area of Feminist Human-Robot Interaction (HRI), the current paper focuses on exploring how robots could be designed to help women to deal with various common challenges, as pictured in Fig. 1.
[…]
@@
Figure 1: Basic concept: robots could be designed to help women related to crime, illness, and day-to-day challenges
@@
As we look toward the future, in which technologies like Artificial Intelligence (AI) and robotics are expected to help people to live better lives, we find similar indications of potential marginalization: For example, reported examples of gender bias in AI include Google preferentially showing high-paying ads to men, Amazon’s hiring system preferring applications from men, Google Translate defaulting to male pronouns, LinkedIn suggesting that female names are mistaken, and Google searches for women often turning up erotic content [2]. As another example, Autonomous Vehicles (AVs) are poised to contribute to safe, efficient transport, but even basic vehicular features such as seat belts, airbags, and crash dummies have up until recently seemingly not taken female body sizes into account, leading to an increased rate of injury in traffic accidents (47% higher between 1998 and 2008 in the US). Furthermore, as Winkle et al. describe, female opinions have also been highly underrepresented in HRI studies, as with Human-Computer Interaction [3]: robot designers are generally not female, and examples of tackling women’s aims and needs appear to be few. (Note: In using the term "women" here, we do not focus on non-binary or gender-fluid cases, although our ideation might also be relevant to some degree for such groups; as well, we follow the typical convention that "sex" refers to biological state and "gender" to social identity.)
One barrier is that the designated design space is so expansive: Women face various challenges that can seem "wicked" and difficult to solve due to their complex, unclear, and interconnected nature. Furthermore, robotic capabilities are ever-expanding; various kinds of robots can be imagined (e.g., drones, wheeled robots, humanoids, exoskeletons, etc.) that could either be carried by a woman, located in an environment, or sent by an authority such as the police or Emergency Medical Services (EMS). Thus, the context suggested the usefulness of following a critical approach that can elicit new insights into both theory and practice.
@@
The remainder of the paper is structured as follows: In Section II we briefly frame and motivate our work in comparison to previous work. Section III describes the speculative prototyping approach we adopted, as well as challenges and opportunities identified. Section IV goes deeper into one specific scenario via a prototype, and Section V discusses results.
@@
  • Theoretical. We identify and explore some new opportunities and challenges related to useful tasks for robot designers to help a large user group (women).
  • Practical. We explore one scenario in more depth, reporting on a new proof-of-concept of a robot.
@@ II RELATED WORK
Previous related work seems to have mostly focused on topics outside of the scope of the current paper, such as differences in how women and men perceive robots, or sex robots. For example, women were observed to place more trust than men in the idea of a security robot [4], and sex robots, rather than promoting objectification of women, could actually foster liberalization and sex-positivity [5]).
@@
However, a few studies have explored how robot designers could address challenges specific to women. One work that is highly relevant to the current topic is by Winkle et al., who proposed Feminist HRI, highlighting the importance of examining and challenging power, in considering diverse perspectives, emotions, embodiment, and "low power" users, to achieve responsible design [6]. As well, Winkle and Mulvihill explored how robots could be used to abuse women in a domestic violence (DV) scenario, toward supporting risk mitigation [7]. The authors comment briefly on possibilities of using ambient sensor data or mobile apps to help detect DV, yet advise caution–referring to a study by Cookson et al. that discusses how digital interventions can be overly hyped and under-deliver, have unintended consequences and hidden costs, and alone are no "magical bullet" [8]. A difference with the current work is that these studies did not seek to speculate on what kinds of specific tasks future robots could one day perform to help women, as we seek to do here. For example, helping could also be possible one day for robots in a DV scenario, as perpetrators typically try to isolate their victims and avoid having their crimes recorded or observed–a task that could be complicated by the close presence of a robot outfitted with sensors.
@@
In contrast to the above studies, much work has looked at how robots could use sensing to help people in general. For example, in our own previous work, we explored how a robot could seek to defend a person who is being attacked [9]. We also experimented with combining object detection in RGB and thermal trace detection to detect activities, which inspired our prototype in the current paper [10]. However, these works have not focused on the unique requirements of women. Thus, it seemed like the question of how robots could help women could benefit from some additional attention.
III METHOD
@@
To identify challenges and opportunities, a speculative prototyping approach was followed. Speculative prototyping aims to "drag" potentially important future scenarios from the foggy realm of imagination closer to the real world, visualizing in a thought-provoking, clear, visceral manner, as a means to begin a process of critical exploration [11]. Thus, rapid ideation sessions were first conducted to identify challenges affecting women. Next, opportunities were identified. Ideas were recorded without judgement, then afterwards merged and reworked with priority attributed to those that seemed valuable, different, and requiring further study. (In doing so, we aimed to roughly align ourselves with Suvin’s concept of "cognitive estrangement", described by Bartolotta in terms of seeking to capture "fictional world-affordances imbued with cognitive potential" [12].) Thus, some representative scenarios were chosen and transformed into sketches. Furthermore, one of these sketches was further developed into a physical prototype, which was used to also gain some insight into practical concerns.
@@ III-A Challenges Faced by Women
What then are some of the challenges that seem to be predominantly faced by women? In considering this question here, we adopt a zemiological perspective; i.e., a broad view that can encompass various potential harms, in relation to health, crime, and day-to-day activities.
@@ III-A1 Crimes
Crimes that disproportionately affect women include voyeurism, sexting without permission, stalking, indecent exposure, domestic abuse (including intimate partner abuse), rape (including date rape, acquaintance rape, marital rape, gang rape, and war rape), molestation (grabbing, groping), child abuse and grooming, forced prostitution (including sex trafficking, sexual slavery, forced marriage, revenge porn, and forced participation in the production of pornography), and other forms of sexual assault, harassment, bullying, and violence [13]. [^4] https://pcar.org/about-sexual-violence/adults We note that these terms are not exclusive but often contain some overlap. As well, crimes can be further detailed in terms of individual actions. For example, physical violence can involve slapping, pushing, hitting, kicking, hair-pulling, choking, burning, or assault with objects such as knives or guns. Also, even such non-sexual violence can relate to sex, since women who are smaller can be seen as easier targets [14].
@@
Such distressing experiences are not limited to any one country. For example, a report from 2002 indicated that two thirds of Swedish women have been sexually harassed or experienced violence [15]: In more detail, over half of women have been sexually harassed, and around half have experienced violence (a quarter physical violence, one in three sexual violence), including around a quarter of young women 18-24 just in the past year; one out of every five women has also been threatened by a man. Numbers are also disturbing in the US, where around two thirds of women have experienced violence, and around one fifth of women have been raped, usually by someone they knew [13]. In some regions, such as West or South Asia and Africa, women can also face risks of female genital mutilation, honor killings, stonings, acid throwing, blinding, persecution due to alleged witchcraft, and forced suicide (e.g. Sati, in which a widow is forced to die on her husband’s funeral pyre [^5] https://en.wikipedia.org/wiki/Sati_(practice)).
@@
The crimes above can be categorized as occurring remotely, or in proximity. Of the first non-contact category, voyeurism, also related to scopophilia, upskirting and peeping, could be one useful problem for initial exploration: In contrast to more serious crimes where the "stakes" can be extremely high if a mistake is made, voyeurism has been called a "harmless wrong", as victims often don’t know they have been victimized and thus do not suffer trauma [16]. Moreover, current legislature regarding filming without consent seems to support technological approaches to limit video voyeurism, and it is also arguably clear how voyeurism can be stopped, by preventing a criminal from viewing and filming victims. As well, voyeurism is an important, widespread problem: Although actual numbers are difficult to predict, voyeurism has been described as an epidemic in Korea, with over 6000 cases reported each year [17]. When footage is spread, the result can be "social death", causing some women to attempt suicide. For such reasons, in 2018, approximately 20,000 women marched to protest spy cameras, with 200,000 signing a petition, which resulted in a government plan to hire 8000 workers to tackle the problem. As well, in Sweden, in an anonymous survey of 2,450 randomly selected 18–60 year-olds conducted in 1996, 7.7% of respondents described deriving sexual satisfaction from spying on others having sex [18]. Thus, it could be useful for women if voyeurism could be more easily detected and prevented.
@@
Of crimes that occur in proximity, two kinds that seem more difficult to tackle, but highly important, include DV and rape. For example, in Sweden, there are typically no witnesses to DV, as the attacks occur "behind closed doors" [19]. As such, various statistics on DV can be found, but the real numbers, and hence the true extent of this problem, are unknown. As well, hospital records are often generated as a result of attacks, but police are usually not allowed to access these data; victims can also feel retraumatized and uncared for when interacting with healthcare staff.
@@
Furthermore, recent discourse has highlighted difficulties in Sweden related to immigration, integration, and crime [20]: Especially victims coming from other cultures can find themselves in a position of weakness, segregated, and like they have no one to talk to: They might not know the local language (Swedish), be allowed to go outside, or be aware of Swedish norms (i.e., no one might explicitly tell them that is not okay for a man to beat a woman). They might also visit religious centers that reinforce non-Swedish views, like that they must obey the men in their families, or that they could be punished or killed for having a local boyfriend. For example, views in Africa on DV can be startlingly different, with acceptance of wife beating at 77% in Mali and Uganda; overall, "51% of African women report that being beaten by their husbands is justified if they either go out without permission, neglect the children, argue back, refuse to have sex, or burn the food." [^6] https://blogs.worldbank.org/en/africacan/domestic-violence-and-poverty-in-africa-when-the-husbands-beating-stick-is-like-butter
Another key problem with DV is that repeat offenses are common [21]. For example, officers might pick up a criminal on Friday for beating a woman, then release him on Saturday, in a pattern that repeats itself each week. The perpetrator as well can escalate, beating harder and harder each time, which can end in murder. Some mechanism is needed to break such damaging spirals.
[…]
@@
Then, what might be needed? As above, one challenge is getting information to the victims, so they are no longer alone. Also, women who are victims need to be protected; something should be done for them without them having to start the process.
@@
Another important problem to combat is rape, which shares many similarities with DV: Rapes are usually conducted by perpetrators known to the victim; the crime is committed in close contact; its repercussions are among the most serious of crimes (leading to chronic health problems such as PTSD, MDD, and substance abuse); help-seeking victims often experience a "second rape" in being repeatedly interrogated, doubted, blamed, warned, and discouraged; and a chronic failure to report or investigate rape has been noted [24] [25]. In some countries, like the US, clearance rates for rape also appear to be low (e.g. in the 60 percent range, compared to Japan in the high 90s, with rapes being about 100 times more common) [26].
@@
Currently, in Sweden, one obstacle to preventing rape through adopting new technologies comes from the legal system. For example, the European Parliament, in its recent "Artificial Intelligence Act", has banned real-time detection, on the basis of the potential for mistakes. [^11] https://www.europarl.europa.eu/news/en/press-room/20240308IPR19015/artificial-intelligence-act-meps-adopt-landmark-law More concretely, in an urban environment with surveillance cameras, it is not permitted to detect rape in real time and alert police. Rape must happen first–everyone is required to wait–then the images can be used, after judicial approval is obtained. As such, currently much of the potential use of AI is lost, and the integrity of victims is not upheld. Thus, here the challenge seems to be not merely technical but also political. While it’s clear that current recognition systems are imperfect, various examples of working real-time systems exist, not just in pacemakers and brakes, but also employing computer vision, from Tesla cars doing real-time inference from cameras, to real-time pixelization of faces on television. Furthermore, a basis exists for how to handle imperfect recognition in standards such as SOTIF (ISO/PAS 21448). [^12] https://www.iso.org/standard/77490.html Thus, an example of a prototype of how robots could detect and prevent crime in real-time at an early stage could also be useful as a way to affect thought regarding legislature.
@@
Another problem involves bottlenecks in processing evidence. Disturbingly, it seems there are hundreds of thousands of untested "rape kits" in the US, some of which are not processed even after 30 years [27]. [^13] https://usafacts.org/articles/how-many-rape-kits-are-awaiting-testing-in-the-us-see-the-data-by-state/ This forensic evidence, which is costly and requires hours of invasive handling of victims, is sometimes not sent onward for analysis or cannot be handled due to insufficient resources. Various potential causes have been put forward such as budget cuts; victim-blaming and bias against women and victims of sex crimes; slow workflows due to fear of contamination and inefficient or redundant testing/auditing requirements; lack of a tracking system for forensic evidence; and misunderstandings of processes surrounding rape kits. In some cases, rape kits have been destroyed before testing without notifying victims. In some other countries, lack of rape kits or trained examiners can also be a problem. The result can include decreased community trust in police, as well as missed opportunities to identify serial perpetrators and prevent new rapes and other crimes–as rapists are often also guilty of other crimes such as burglary, assault, and murder.

### v2

@@ Abstract
Robots are being designed to help people in an increasing variety of settings–but seemingly little attention has been given so far to the specific needs of women, who represent roughly half of the world’s population but are underrepresented in robotics. Here we used a speculative prototyping approach to explore this expansive design space: First, we identified some challenges that disproportionately affect women in relation to crime, health, and daily activities, as well as opportunities for designers, which were visualized in five sketches. Then, one of the sketched scenarios was further explored by developing a prototype, of a drone equipped with computer vision to detect hidden cameras that could be used to spy on women. While object detection introduced some errors, hidden cameras were identified with a reasonable accuracy of 80% (Intersection over Union (IoU) score: 0.40). Our aim is that these results could help spark discussion and inspire designers, toward realizing a safer, more inclusive future through responsible use of technology.
@@ I INTRODUCTION
Within the area of feminist Human-Robot Interaction (HRI), the current paper explores how robots could be designed to help women to deal with various common challenges, as pictured in Fig. 1.
Discrepancies can sometimes be observed between how we would like the world to function, and how the world actually functions: We believe that people should be treated equally, with similar rights and opportunities, and that women are an important group to consider, in line with the concept of gender mainstreaming and the UN’s Sustainable Development Goal 5.111eige.europa.eu/gender-mainstreaming/what-is-gender-mainstreaming,222unwomen.org/en/node/36060 Yet despite constituting roughly half of the human population, women have been historically marginalized, underrepresented, ignored, and restricted [1].333womansday.com/life/real-women/a55991/no-women-allowed Some beliefs can seem humorous in retrospect, like that women might not be able to ride trains as their uteruses might fly out of their bodies due to high speeds. Other more sobering examples question rights that might seem fundamental, like the rights to vote, run, or defend, which have been granted only recently in some countries (e.g., the right to vote was granted in Liechtenstein only in 1984, the right to run with men in the Boston Marathon in the United States (US) only in 1972, and the right to carry out some combat jobs in the US army only in 2016).
@@
Figure 1: Basic concept: robots could be designed to help women with challenges related to crime, health, and daily activities
@@
As we look toward the future, in which technologies like Artificial Intelligence (AI) and robotics are expected to help people to live better lives, we find similar indications of potential marginalization: For example, reported examples of gender bias in AI include Google preferentially showing high-paying ads to men, Google Translate defaulting to male pronouns, Amazon’s hiring system preferring applications from men, and LinkedIn suggesting that female names are mistaken [2]. [^4] https://www.reuters.com/article/idUSKCN1MK0AG/  [^5] https://qz.com/775597/linkedins-lnkd-search-algorithm-apparently-favored-men-until-this-week As another example, Autonomous Vehicles (AVs) are poised to contribute to safe, efficient transport, but even basic automotive features such as seat belts, airbags, and crash dummies have up until recently seemingly not taken female body sizes into account, leading to an increased rate of injury in traffic accidents (47% higher between 1998 and 2008 in the US); as well, cars’ speech recognition systems have had trouble understanding female voices [3]. Furthermore, as Winkle et al. describe, female opinions have been highly underrepresented in HRI studies, as with Human-Computer Interaction [4]: robot designers are generally not female, and examples of tackling women’s needs appear to be scarce. (Note: In using the term "women" here, we do not focus on non-binary or gender-fluid cases, although our ideation might also be relevant to some degree for such groups; as well, we follow the typical convention that "sex" refers to biological state and "gender" to social identity.)
One barrier is that the designated design space is expansive: Women face various challenges that can seem "wicked", complex, unclear, interwoven, and difficult to solve; robots also can have various capabilities, forms (e.g., flying, wheeled, humanoid, or exoskeletal), and deployment configurations (e.g., carried, located nearby, or sent from a police station or hospital), etc. Thus, a rough understanding of the "big picture" would seem useful for designers and policy makers to select meaningful challenges to tackle, consider potential solutions, and make informed decisions. Here, we follow a critical speculative prototyping approach intended to elicit new insights into both theory and practice.
@@
The remainder of the paper is structured as follows: In Section II, we briefly frame and motivate our work in comparison to previous work. Section III describes how we identified challenges and opportunities. Section IV goes deeper into one specific scenario via a prototype, and Section V discusses results.
@@
  • Theoretical. We identify and explore some new challenges and opportunities related to useful tasks for robot designers to help a large user group (women).
  • Practical. We explore one scenario in more depth, reporting on a new proof-of-concept for a robot.
@@ II RELATED WORK
Previous work related to women and HRI seems to have mostly focused on topics outside of the scope of the current paper, such as differences in how women and men perceive robots, or sex robots. For example, women were observed to place more trust than men in the idea of a security robot [5], and sex robots, rather than promoting objectification of women, could actually foster liberalization and sex-positivity [6]).
@@
However, a few studies have explored how robot designers could address challenges specific to women. One work that is highly relevant to the current topic is by Winkle et al., who proposed feminist HRI–highlighting the importance of examining and challenging power, in considering diverse perspectives, emotions, embodiments, and "low power" users, to achieve responsible design [7]. As well, Winkle and Mulvihill explored how robots could be used to abuse women in a domestic violence (DV) scenario, toward mitigating risks [8]. The authors comment briefly on possibilities of using ambient sensor data or mobile apps to help detect DV, yet advise caution–referring to a study by Cookson et al. that discusses how digital interventions can be overly hyped and under-deliver, have unintended consequences and hidden costs, and alone are no "magical bullet" [9]. A difference with the current work is that these studies did not seek to speculate about what kinds of tasks future robots could one day perform to help women. For example, the presence of robots with sensors could also one day help in a DV scenario, by complicating perpetrators’ efforts to isolate victims and avoid having crimes recorded or observed.
@@
Much work has also looked at how robots could use sensing to help people in general. For example, we previously explored how a robot could seek to defend a person who is being attacked [10]. We also experimented with combining object detection and thermal trace detection to infer activities, which inspired our prototype in the current paper [11]. However, these works did not focus on the unique requirements of women. Thus, it seemed useful to further explore the question of how robots could help women.
III SPECULATION
@@
To provide both theoretical and practical insights, a speculative prototyping approach was followed. Speculative prototyping aims to "drag" potentially important future scenarios from the foggy realm of imagination closer to the real world, visualizing in a thought-provoking manner, as a means of embarking on a process of critical exploration [12]. As a first step, rapid ideation sessions were conducted among the authors, who contained a female member and had some experience with robotics, AI, computer vision, criminology, and health technology. For both challenges and opportunities, ideas were recorded without judgement, then afterwards merged, refined by surveying related literature, and reworked with priority attributed to those that seemed valuable, different, and requiring further study. In doing so, we aimed to roughly align ourselves with Suvin’s concept of "cognitive estrangement", described by Bartolotta in terms of seeking to capture "fictional world-affordances imbued with cognitive potential" [13]. We also adopted a zemiological perspective [8]; i.e., a broad view that encompasses various potential harms, in relation to three topics we identified: crime, health, and daily activities.
@@ III-A1 Crimes
Crimes that disproportionately affect women can occur globally or locally, and at distance or in proximity. In general, crimes that don’t require close contact include voyeurism, sexting without permission, stalking, and indecent exposure; crimes that require contact include DV (including intimate partner abuse), rape (including date, acquaintance, marital, and gang rapes), molestation (grabbing, groping), child abuse and grooming, forced prostitution (including sex trafficking, sexual slavery, forced marriage, revenge porn, and forced participation in the production of pornography), and other forms of sexual assault, harassment, bullying, and violence [14]. [^6] https://pcar.org/about-sexual-violence/adults In some regions of Asia and Africa, women can also face risks of female genital mutilation, honor killings, stonings, acid throwing, blinding, persecution due to alleged witchcraft, war rape, and forced suicide (e.g., Sati, in which a widow is forced to die on her husband’s funeral pyre [^7] https://en.wikipedia.org/wiki/Sati_(practice)).
Various complexities exist: For example, some of the above terms overlap, and even non-sexual violence can relate to sex, since women who are smaller can be seen as easier targets [15]. As well, crimes can be further detailed in terms of individual actions; e.g., physical violence can involve slapping, pushing, hitting, kicking, hair-pulling, choking, burning, or assault with objects such as knives or guns.
@@
Yet despite these complexities, crimes would be important to mitigate since they affect many women. For example, a report from 2002 indicated that two thirds of Swedish women have been sexually harassed or experienced violence [16]: Over half of women have been sexually harassed, and around half have experienced violence (a quarter physical violence, one in three sexual violence), including around a quarter of young women 18-24 just in the past year; one out of every five women has also been threatened by a man. Numbers are also disturbing in the US, where around two thirds of women have experienced violence, and around one fifth of women have been raped, usually by someone they knew [14].
@@
From the large pool of crimes described, three crimes were selected to explore in greater detail, one remote, and two requiring contact. From the remote category, voyeurism, also related to scopophilia, upskirting and peeping, seemed useful to explore, since it might be relatively easy to begin with, and is an important, widespread problem: The cost of a failed intervention could be lower than for more serious crimes, given that victims often don’t know they have been victimized and thus don’t suffer trauma [17]. Moreover, current legislature regarding filming without consent seems to support technological approaches to limit video voyeurism, and it seems relatively clear how voyeurism can be stopped, by preventing a criminal from viewing and filming victims. Furthermore, although actual numbers are difficult to predict, voyeurism has been described as an epidemic in Korea, with over 6000 cases reported each year [18]. When footage is spread, the result can be "social death", causing some women to attempt suicide. For such reasons, in 2018, approximately 20,000 women marched to protest spy cameras, with 200,000 signing a petition, which resulted in a government plan to hire 8000 workers to tackle the problem. As well, in Sweden, in an anonymous survey of 2,450 randomly selected 18–60 year-olds conducted in 1996, 7.7% of respondents described deriving sexual satisfaction from spying on others having sex [19]. Thus, it could be useful for women if voyeurism could be more easily detected and prevented.
@@
Of crimes that require proximity, two kinds that seem more difficult to tackle, but highly important, include DV and rape. For example, in Sweden, there are typically no witnesses to DV, as the attacks occur "behind closed doors" [20]. As such, various statistics on DV can be found, but the real numbers, and hence the true extent of this problem, are unknown. As well, hospital records are often generated as a result of attacks, but police are usually not allowed to access these data; victims can also feel retraumatized and uncared for when interacting with healthcare staff.
@@
Furthermore, recent discourse has highlighted difficulties in Sweden related to immigration and integration [21]: Especially victims coming from other cultures can find themselves in a position of weakness–segregated, and with no one to talk to: They might not be capable of communicating in the local language, allowed to go outside, or aware of Swedish norms (i.e., no one might explicitly tell them that is not okay for a man to beat a woman). They might also visit religious centers that reinforce non-Swedish views, like that they must obey the men in their families, or that they could be punished or killed for having a local boyfriend. For example, views in Africa on DV can be startlingly different, with acceptance of wife-beating at 77% in Mali and Uganda; overall, "51% of African women report that being beaten by their husbands is justified if they either go out without permission, neglect the children, argue back, refuse to have sex, or burn the food." [^8] https://blogs.worldbank.org/en/africacan/domestic-violence-and-poverty-in-africa-when-the-husbands-beating-stick-is-like-butter Another key problem with DV is that repeat offenses are common over the long term [22]. For example, officers might pick up a criminal on Friday for beating a woman, then release him on Saturday, in a pattern that repeats itself each week. The perpetrator as well can escalate, beating harder and harder each time, which can end in murder. Some mechanism is needed to break such damaging spirals.
Various suggestions have been made: For example, phone calls to DV helplines can be masked, and women freed from family phone plans on request. [^9] https://ksltv.com/602266/fcc-adopts-new-cellphone-rules-designed-to-keep-domestic-violence-victims-safer As well, help can be sought by drawing a black dot on one’s hand, selecting a red pen at a hospital, or uttering keywords like "Angela" or "Minotaur" at a bar.101010www.bbc.com/news/blogs-trending-34326137,111111themighty.com/2020/01/domestic-violence-preventionsign- red-marker,121212canadianwomen.org/signal-for-help However, it’s not clear how effective such existing strategies are: Backlash effects could exist; if an aggressor learns about such a communication, the victim could be beaten more. Furthermore, perpetrators typically seek to "gaslight" or manipulate victims into doubting themselves, by establishing narratives to maintain power and control–e.g., telling victims that they are crazy, useless, or to blame, and that this is how things must be [23]. As a result of repeatedly experiencing traumatic and uncontrollable events, including physical and psychological abuse, victims often have little self-esteem left, and frequently develop posttraumatic stress disorder (PTSD), major depressive disorder (MDD), or "learned helplessness"–which contribute to submissiveness and reduce a woman’s belief that her actions could lead to a positive result [24]. Thus, while successful arrests could help to reduce repeated offences by approximately half [22], as with other crimes such as sex trafficking or forced marriage, battered women often feel trapped, and it’s hard for women themselves to break free.
@@
Then, what might be needed? As above, one challenge is getting information to the victims, so they are no longer alone. Also, women who are victims need to be protected; something should be done without them having to start the process.
@@
Another important problem to combat is rape, which shares similarities with DV: Rapes are usually conducted by perpetrators known to the victim; the crime is committed in close contact; its repercussions are among the most serious of crimes (leading to chronic health problems such as PTSD, MDD, and substance abuse); help-seeking victims often experience a "second rape" in being repeatedly interrogated, doubted, blamed, warned, and discouraged; and a chronic failure to report or investigate rape has been noted [25] [26]. In some countries, like the US, clearance rates for rape also appear to be low (e.g., in the 60 percent range, compared to Japan in the high 90s, with rapes being about 100 times more common) [27].
@@
Challenges include legislation and processing of evidence. For example, the European Parliament, in its recent "Artificial Intelligence Act", has banned real-time detection, on the basis of the potential for mistakes.131313europarl.europa.eu/news/en/press-room/20240308IPR19015/artificial-intelligence-act-meps-adopt-landmark-law More concretely, in an urban environment with surveillance cameras, it is not permitted to detect rape in real time and alert police. Rape must happen first–everyone is required to wait–then images can be used, after judicial approval is obtained. As such, currently much of the potential use of AI is lost, and the integrity of victims is not upheld. While it’s clear that current recognition systems are imperfect, various examples of working real-time systems exist, not just in pacemakers and brakes, but also employing computer vision, from Tesla cars doing real-time inference from cameras, to real-time pixelization of faces on television. Furthermore, a basis exists for how to handle imperfect recognition in standards such as SOTIF (ISO/PAS 21448). [^14] https://www.iso.org/standard/77490.html Thus, here the challenge seems to be not merely technical but also political; a prototyping example of how robots could detect and prevent crime in real-time at an early stage could also be useful as a way to affect thought regarding legislature.
@@
Another problem involves bottlenecks in processing evidence. Disturbingly, it seems there are hundreds of thousands of untested "rape kits" in the US, some of which are not processed even after 30 years [28]. [^15] https://usafacts.org/articles/how-many-rape-kits-are-awaiting-testing-in-the-us-see-the-data-by-state This forensic evidence, which is costly and requires hours of invasive handling of victims, is sometimes not sent onward for analysis or cannot be handled. Various potential causes have been put forward such as budget cuts; victim-blaming and bias against women and victims of sex crimes; slow workflows due to fear of contamination, and inefficient or redundant testing/auditing requirements; lack of a tracking system for forensic evidence; and misunderstandings of processes to follow. In some cases, rape kits have been destroyed before testing without notifying victims, and in some areas of the world, lack of rape kits or trained examiners can also be a problem. The result can include decreased community trust in police, as well as missed opportunities to identify serial perpetrators and prevent new rapes and other crimes–as rapists are often also guilty of other crimes such as burglary, assault, and murder.